Determine the TWA of the following chlorine exposure. Is the worker overexposed to the ACGIH TLV updated in 2017?

2 hours at None detected

1 hour at 0.2 ppm

3 hours at 0.1 ppm

2 hours at 0.15 ppm
Answer #1

TWA or the time weighted average is the average exposure to a harmful substance for a specific period.

The TWA can be calculated by:

TWA= Citi + C2t2 + ... + Cntn ti + t2 + ... + tn.

where C_1,C_2,C_3,... are the concentration of the substance and t_1,t_2,t_3,... are the time of exposure to the harmful substance.

In the given problem, the employee spends 2 hours where there is no chlorine detected

So C_1=0 and t_1=2 .

The employee also spends 1 hour at 0.2 ppm

So C_2=0.2 .  .t_2=1

He also spends another 3 hours at 0.1 ppm

i.e. C_3=0.1 .  .t_3=3

and also spends another 2 hours at 0.15 ppm

C_4=0.15 .  .t_4=2

Therefore the employee’s eight-hour TWA is:

TWA=\frac{\left (0\times 2 \right )+\left (0.2\times 1 \right )+\left (0.1\times 3 \right )+\left (0.15\times 2 \right )}{8}

i.e.  .TWA=\frac{0+0.2+0.3+0.3}{8}

TWA=\frac{0.8}{8}=0.1

Therefore employee's 8 hour TWA of chlorine exoposure is 0.1 ppm.

Now, the ACGIH TLV for Chlorine is 0.5 ppm. So, we can say that the worker's TWA (0.1 ppm) is less than that of the TLV value for chlorine (0.5 ppm).

Therefore the worker is not overexposed to chlorine.